# Generalized and Efficient 2D Gaussian Splatting for Arbitrary-scale Super-Resolution

### v2

@@ 3.2 GSASR: Architecture and Modules
where $Q$ is from Gaussian embedding $\mathbf{E}$, $K$ and $V$ are from image feature $\mathbf{F}$. The $i$-th embedding $\mathbf{E}_{i}$ at position $p_{i}$ interacts with the $j$-th pixel features within the window. $B$ is the relative position bias in the feature map.
Table 2: Quantitative comparison between representative ASR models and our GSASR. All models use the same RDN [80] backbone as the feature extraction encoder. We test on DIV2K and LSDIR [36] datasets [55] with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$, ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$. The best results are highlighted in red. The PSNR and SSIM indices are computed in the Y channel of Ycbcr space.
Scale Metrics Backbone: RDN
Testing Dataset: DIV2K Testing Dataset: LSDIR
Meta
-SR
LIIF LTE SRNO LINF LMF
Ciao
-SR
Gaussian
-SR
GSASR
Meta
-SR
LIIF LTE SRNO LINF LMF
Ciao
-SR
Gaussian
-SR
GSASR
$\times$2 PSNR 36.54 36.37 36.41 36.53 36.53 36.48 36.67 36.54 36.73 31.97 31.87 31.96 32.10 31.94 32.02 32.16 31.95 32.26
SSIM 0.9483 0.9481 0.9484 0.9493 0.9483 0.9487 0.9494 0.9484 0.9500 0.9224 0.9221 0.9228 0.9247 0.9223 0.9234 0.9247 0.9228 0.9261
LPIPS 0.0816 0.0828 0.0816 0.0783 0.0821 0.0806 0.0797 0.0829 0.0757 0.0872 0.0876 0.0861 0.0832 0.0874 0.0866 0.0839 0.0875 0.0812
DISTS 0.0542 0.0543 0.0535 0.0528 0.0541 0.0540 0.0524 0.0543 0.0509 0.0644 0.0645 0.0635 0.0623 0.0645 0.0639 0.0615 0.0640 0.0603
$\times$3 PSNR 32.79 32.68 32.73 32.83 32.77 32.78 32.89 32.74 32.97 28.34 28.26 28.34 28.45 28.30 28.37 28.46 28.33 28.56
SSIM 0.8935 0.8934 0.8938 0.8951 0.8936 0.8947 0.8952 0.8939 0.8970 0.8399 0.8398 0.8409 0.8435 0.8398 0.8421 0.8432 0.8407 0.8462
LPIPS 0.1850 0.1842 0.1837 0.1817 0.1846 0.1807 0.1821 0.1847 0.1768 0.1997 0.1983 0.1979 0.1947 0.1995 0.1954 0.1953 0.1993 0.1908
DISTS 0.1000 0.0998 0.0987 0.0980 0.0999 0.0984 0.0965 0.0998 0.0955 0.1192 0.1188 0.1176 0.1159 0.1194 0.1169 0.1139 0.1189 0.1139
$\times$4 PSNR 30.78 30.71 30.75 30.85 30.77 30.81 30.91 30.76 30.96 26.53 26.48 26.54 26.64 26.51 26.58 26.66 26.53 26.73
SSIM 0.8455 0.8449 0.8459 0.8478 0.8453 0.8467 0.8481 0.8457 0.8500 0.7724 0.7714 0.7734 0.7767 0.7719 0.7744 0.7770 0.7727 0.7801
LPIPS 0.2568 0.2566 0.2558 0.2525 0.2574 0.2546 0.2525 0.2570 0.2505 0.2831 0.2838 0.2822 0.2771 0.2840 0.2817 0.2768 0.2837 0.2752
DISTS 0.1341 0.1354 0.1341 0.1330 0.1355 0.1341 0.1327 0.1347 0.1288 0.1581 0.1603 0.1589 0.1567 0.1609 0.1584 0.1563 0.1595 0.1533
$\times$6 PSNR 28.40 28.44 28.49 28.57 28.48 28.52 28.61 28.46 28.65 24.55 24.59 24.64 24.70 24.61 24.66 24.73 24.60 24.77
SSIM 0.7715 0.7739 0.7749 0.7772 0.7739 0.7757 0.7777 0.7727 0.7800 0.6756 0.6791 0.6806 0.6841 0.6787 0.6815 0.6853 0.6772 0.6881
LPIPS 0.3505 0.3468 0.3551 0.3510 0.3523 0.3534 0.3452 0.3584 0.3443 0.3960 0.3915 0.4020 0.3964 0.3961 0.4009 0.3864 0.4043 0.3867
DISTS 0.1837 0.1882 0.1878 0.1861 0.1890 0.1873 0.1867 0.1880 0.1817 0.2110 0.2166 0.2161 0.2131 0.2181 0.2154 0.2131 0.2167 0.2093
$\times$8 PSNR 26.97 27.07 27.13 27.20 27.11 27.16 27.24 26.99 27.28 23.44 23.49 23.54 23.60 23.52 23.57 23.63 23.45 23.63
SSIM 0.7222 0.7274 0.7286 0.7308 0.7272 0.7293 0.7317 0.7228 0.7339 0.6151 0.6221 0.6230 0.6264 0.6211 0.6240 0.6281 0.6156 0.6299
LPIPS 0.4133 0.4072 0.4214 0.4166 0.4163 0.4188 0.4075 0.4290 0.4060 0.4679 0.4627 0.4798 0.4729 0.4711 0.4772 0.4592 0.4875 0.4593
DISTS 0.2181 0.2251 0.2268 0.2248 0.2273 0.2260 0.2244 0.2285 0.2192 0.2489 0.2562 0.2569 0.2537 0.2588 0.2563 0.2528 0.2601 0.2490
$\times$12 PSNR 25.23 25.39 25.45 25.52 25.41 25.48 25.55 25.12 25.56 22.13 22.21 22.26 22.31 22.22 22.28 22.33 22.06 22.30
SSIM 0.6653 0.6739 0.6748 0.6768 0.6730 0.6754 0.6783 0.6648 0.6794 0.5494 0.5597 0.5605 0.5629 0.5576 0.5608 0.5651 0.5484 0.5656
LPIPS 0.5001 0.4981 0.5180 0.5132 0.5120 0.5153 0.4982 0.5455 0.4947 0.5671 0.5685 0.5899 0.5823 0.5803 0.5884 0.5635 0.6174 0.5621
DISTS 0.2703 0.2809 0.2857 0.2835 0.2855 0.2844 0.2811 0.2921 0.2759 0.3055 0.3138 0.3165 0.3133 0.3187 0.3160 0.3116 0.3252 0.3077
$\times$16 PSNR 24.12 24.30 24.36 24.42 24.30 24.38 24.44 23.93 24.44 19.34 21.40 21.45 21.49 21.41 21.45 21.50 21.20 21.47
SSIM 0.6346 0.6446 0.6452 0.6467 0.6433 0.6456 0.6484 0.6340 0.6487 0.4813 0.5273 0.5277 0.5294 0.5253 0.5269 0.5318 0.5156 0.5318
LPIPS 0.5606 0.5580 0.5788 0.5739 0.5751 0.5760 0.5572 0.6166 0.5537 0.6762 0.6343 0.6556 0.6487 0.6488 0.6598 0.6281 0.6932 0.6261
DISTS 0.3117 0.3243 0.3312 0.3288 0.3324 0.3295 0.3268 0.3455 0.3211 0.3845 0.3566 0.3605 0.3575 0.3638 0.3596 0.3562 0.3745 0.3512
$\times$18 PSNR 23.69 23.87 23.93 23.99 23.87 23.96 24.01 23.48 23.98 20.99 21.09 21.14 21.17 21.10 21.13 21.18 20.88 21.14
SSIM 0.6244 0.6346 0.6349 0.6362 0.6333 0.6353 0.6380 0.6239 0.6379 0.5056 0.5165 0.5168 0.5181 0.5147 0.5160 0.5205 0.5056 0.5203
LPIPS 0.5776 0.5805 0.6013 0.5968 0.5957 0.5987 0.5796 0.6418 0.5766 0.6489 0.6575 0.6490 0.6720 0.6696 0.6822 0.6511 0.7188 0.6497
DISTS 0.3303 0.3419 0.3499 0.3476 0.3513 0.3480 0.3455 0.3676 0.3391 0.3633 0.3733 0.3784 0.3752 0.3814 0.3764 0.3740 0.3944 0.3682
$\times$24 PSNR 22.71 22.87 22.92 22.97 22.87 22.93 22.99 22.49 22.96 19.26 20.39 20.44 20.47 20.41 20.43 20.48 20.16 20.42
SSIM 0.6041 0.6139 0.6138 0.6145 0.6128 0.6136 0.6166 0.6048 0.6162 0.4728 0.4967 0.4965 0.4973 0.4953 0.4955 0.4995 0.4873 0.4985
LPIPS 0.6273 0.6301 0.6505 0.6461 0.6423 0.6487 0.6292 0.6939 0.6275 0.7112 0.7073 0.7274 0.7207 0.7143 0.7316 0.7005 0.7701 0.7006
DISTS 0.3830 0.3865 0.3978 0.3955 0.4008 0.3961 0.3932 0.4224 0.3851 0.4205 0.4145 0.4220 0.4185 0.4260 0.4212 0.4172 0.4434 0.4095
$\times$30 PSNR 17.25 22.18 22.23 22.27 32.18 22.18 22.28 21.81 22.23 19.78 19.86 19.91 19.93 19.89 19.90 19.94 19.64 19.86
SSIM 0.5546 0.6016 0.6017 0.6024 0.6010 0.5995 0.6041 0.5947 0.6032 0.4758 0.4842 0.4840 0.4844 0.4835 0.4837 0.4861 0.4770 0.4851
LPIPS 0.6817 0.6652 0.6833 0.6799 0.6757 0.7009 0.6627 0.7247 0.6633 0.7164 0.7418 0.7598 0.7540 0.7468 0.7570 0.7269 0.8009 0.7358
DISTS 0.4405 0.4215 0.4359 0.4332 0.4394 0.4361 0.4301 0.4650 0.4213 0.4509 0.4454 0.4562 0.4520 0.4605 0.4536 0.4495 0.4823 0.4407
@@
Gaussian Interaction Block. In the condition injection block, each Gaussian embedding $\mathbf{E}_{i}$ is processed independently, resulting in a lack of information exchange among different Gaussian embeddings. To address this issue, we stack $L$ Gaussian interaction blocks to enhance the interaction among Gaussian embeddings. The Gaussian interaction block is designed based on Swin-transformer architecture, which iteratively stacks the window attention layers and shifted window attention layers. In attention layers, we perform self-attention for embeddings whose reference positions are located in the same window, similar to the condition injection block. Besides, we introduce a cross-attention layer to make embeddings scale-aware so that the Gaussian embeddings can learn to adjust the properties according to the given scale factor.
@@
Gaussian Primary Head. We employ several disentangled property heads to convert Gaussian embeddings $\mathbf{E}$ into 2D Gaussians $\mathcal{G}$. The $i$-th Gaussian embedding $\mathbf{E}_{i}$ is fed into a lightweight MLP to output its opacity $\alpha_{i}$, color $c_{i}$, stand deviation $\sigma_{i}$, reference offset $o_{i}$, and correlation $\rho_{i}$. Note that the position is obtained by adding the reference position and predicted offset $\mu_{i}=p_{i}+o_{i}$. The five property heads are decoupled so that different heads can be adapted to predict specific properties. We apply the tanh function to ensure $\rho\in[-1,1]$, and use the sigmoid function for $\{\alpha,c,\sigma\}$ to ensure that their values are physically meaningful. No activation function is applied to $o$ since the positions of Gaussians in 2D images are expected to be free.
@@
Rendering. In Eq. 4, the rasterization determines the sampling interval of the 2D Gaussians based on the scale factor. As illustrated in Figure 3, larger upsampling factors use smaller sampling intervals, while the smaller upsampling factors use larger intervals. Due to the excellent representational capacity of Gaussians, high-quality sampling results can be achieved even in areas with complex textures. After rendering the SR image from rasterization, it is compared with the HR ground truth to calculate the loss $L_{1}$.
4 Experiments
4.1 Experimental Setup
Training Details. We utilize the widely-used DIV2K [55] dataset as training set. We follow [10, 2] to fix LR patches to $48\times 48$. The scaling factor $s$ is randomly selected from $[1.0,4.0]$. We firstly crop $48s\times 48s$ GT patch from the original full-size GT image, and then apply the bicubic down-sampling [2] to obtain the corresponding LR patches. We employ EDSR [40] and RDN [80] as the image encoder backbones to extract feature from the input LR images. We train our GSARS on $4$ NVIDIA A100 GPUs for $500,000$ iterations, with a batch size of 64. The initial learning rate is set to $2e^{-4}$, and halves at $250,000$, $400,000$, $450,000$, $475,000$ iterations. The Adam [29] optimizer is utilized. To speed up the convergence of the training progress, we set the number of warm-up iterations to $2,000$.
Table 3: Comparison of computational costs. We report the results using EDSR-backbone [40] as image encoder. Apart from the PSNR/SSIM/LPIPS/DISTS metrics, we report the average inference time (ms) and the GPU memory usage (MB). The best results are highlighted in red.
Scale Computational Cost and Performance Backbone: EDSR
@@ 4.1 Experimental Setup
Comparison Methods. We compare our proposed GSASR with Meta-SR [24], LIIF [10], LTE [32], SRNO [64], LINF [69], CiaoSR [2], LMF [22], and GaussianSR [23].
Evaluation Protocols. We evaluate the competing models on Set5 [1], Set14 [72], DIV2K100 [55], Urban100 [25], BSDS100 [45], Manga109 [46], General100 [16], and LSDIR [36] datasets under scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$. To validate the generalization on higher scaling factors, we further report the results with ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$ scaling factors on DIV2K100 [55] and LSDIR [36] datasets. For performance measurement, we report PSNR, SSIM [63], LPIPS [75], and DISTS [14] indices. The PSNR and SSIM are computed in the Y channel of Ycbcr space. For the comparison of computational cost, we crop $100$ GT patches with $720\times 720$ resolution from DIV2K [55], and use bicubic interpolation to generate the corresponding LR images with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$. We report the average inference time (ms) and GPU memory usage (MB) on a single NVIDIA A100 GPU.
@@ 4.2 Experiment Results
Quantitative Results. Table 1 and Table 2 show the numerical results of our proposed GSASR and other comparison methods. One could see that GSASR outperforms existing methods in terms of both fidelity metrics (PSNR and SSIM [63]) and perceptual quality metrics (LPIPS [75] and DISTS [14]) under both the EDSR [40] and RDN [80] backbones and scaling factors from ×$\times 2$ to ×$\times 30$. It obtains significant improvements over the methods based on INR, exhibiting greater potential for ASR. Compared with the recent GaussianSR method [23], which also employs GS for feature representation, GSASR shows clear advantages. GSASR predicts the position of Gaussians via an end-to-end manner, while GaussianSR [23] simply replaces the position parameters with the fixed position of RGB pixels of the LR image and loses representation capability. Besides, GaussianSR lacks GPU/CUDA-based rasterization to accelerate the inference speed and it fixes the number of Gaussians to $100$, no matter how large the scaling factor is. Therefore, the performance of GaussianSR drops a lot under high magnification factors. On the contrary, we present an efficient 2D GPU/CUDA-based rasterization to embed much more Gaussians to promote ASR performance. As a result, although GSASR is trained with a scaling factor from $1.0$ to $4.0$, it could not only obtain excellent performance for in-distribution scaling factors (i.e., scaling factor is less than $4$), but also show strong capability in dealing with out-of-distribution situations (i.e., the scaling factor is larger than $4$). More results on Set5 [1], Set14 [72], Urban100 [25], BSDS100 [45], Manga109 [46], General100 [16] could be found in the Appendix.
@@
Qualitative Results. Figure 4 and Figure 5 show the qualitative comparison of competing methods. We can see that the compared methods mostly produce blurry details, for example, the holes in Figure 4, while GSASR generates much clearer edges and details, owing to the strong representation ability from 2D Gaussians. More visualization results can be found in the Appendix. To explain why GSASR works well for the ASR task, we visualize the positions together with the shape of Gaussians in Figure 6. One could see that the positions of Gaussians tend to cluster in areas with complex textures (such as the windows), while they are uniformly distributed in flat areas (such as the sky). This validates that our proposed strategy is highly suitable for restoring complex texture details. Meanwhile, from the blue box and green box in Figure 6, one could find that the Gaussians will adaptively adjust their orientations and scales based on the direction and shape of the textures.
@@
Computational Costs. Table 3 shows that GSASR surpasses all the competing methods in most of the fidelity (PSNR, SSIM) and perceptual (LPIPS, DISTS) measures. In terms of speed, GSASR is nearly twice as fast as the state-of-the-art method CiaoSR [2] under ×3,×4$\times$ 3$\times$ 4× 3 , × 4 scaling factors, and speed advantage becomes more significant when handling with larger scaling factors. Thanks to the efficient 2D GPU/CUDA-based rasterization, GSASR could render an image with high speed without sacrificing the reconstruction fidelity. The computational cost with RDN backbone could be found in the Appendix.
@@
Ablation Study. We conduct ablation studies on the number of Gaussians $N$, and the functionality of the reference position. Details are in the Appendix.

### v3

@@ 3.2 GSASR: Architecture and Modules
where $Q$ is from the Gaussian embeddings $\mathbf{E}$, $K$ and $V$ are from image features $\mathbf{F}$. The $i$-th embedding $\mathbf{E}_{i}$ at position $p_{i}$ interacts with the $j$-th pixel feature within the window. $B$ is inherited from the Swin Transformer [34], serving as a learnable position encoding to indicate the relative position between Gaussian embeddings and LR features.
@@
Gaussian Interaction Block. In the condition injection block, each Gaussian embedding $\mathbf{E}_{i}$ is processed independently, resulting in a lack of information exchange among different Gaussian embeddings. To address this issue, we stack $L$ Gaussian interaction blocks to enhance the interaction among Gaussian embeddings. The Gaussian interaction block is designed based on the Swin Transformer architecture [34], which iteratively stacks the window attention layers and shifted window attention layers. In attention layers, we perform self-attention for embeddings whose reference positions are located in the same window, similar to the condition injection block. Besides, we introduce a cross-attention layer to make embeddings scale-aware so that the Gaussian embeddings can learn to adjust the properties according to the given scale factor.
@@
Gaussian Primary Head. We employ five disentangled heads to convert Gaussian embeddings (output from Gaussian Interaction Block) $\mathbf{E}$ into the corresponding five properties of 2D Gaussians $\mathcal{G}$. To be more specific, the $i$-th Gaussian embedding $\mathbf{E}_{i}$ is fed into five lightweight MLP blocks, each of which contains three linear layers and two ReLU activation functions, to output its opacity $\alpha_{i}$, color $c_{i}$, stand deviation $\sigma_{i}$, reference offset $o_{i}$, and correlation $\rho_{i}$. Note that the position is obtained by adding the reference position and predicted offset $\mu_{i}=p_{i}+o_{i}$. The five property heads are decoupled so that different heads can be adapted to predict specific properties. We apply the tanh function to ensure $\rho\in[-1,1]$, and use the sigmoid function for $\{\alpha,c,\sigma\}$ to ensure that their values are physically meaningful. No activation function is applied to $o$ since the positions of Gaussians in 2D images are expected to be free. Finally, the 2D Gaussians $\mathcal{G}$ are calculated from Eq. 1, Eq. 2 with the predicted five properties $\{\alpha,\mu,c,\sigma,\rho\}$.
Table 2: Comparison of computational costs. We report the results using EDSR-backbone [32] as image encoder. Apart from the PSNR/SSIM/LPIPS/DISTS metrics, we also report the average inference time (ms) and the GPU memory usage (MB). PSNR/SSIM are calculated on the Y channel of Ycbcr space. The inference time/GPU memory cost is computed for the whole SR pipeline, including the encoder, decoder and rendering parts. The best results are highlighted in red.
Scale Computational Cost and Performance Backbone: EDSR-baseline
@@
Rendering. In Eq. 4, during rasterization, the scaling factor $s$ serves as a condition to determine the sampling interval in 2D Gaussians $\mathcal{G}$. As illustrated in Fig. 3, since the rasterization is parameter-free, for different scaling factors $s$, it will adaptively adjust the sampling density to render images with different sizes. Larger scaling factors lead to smaller sampling intervals, vice versa. Therefore, GSASR is equipped with powerful generalization capability to arbitrary scaling factors, which are even unseen in the training process. In the training stage, the rendered SR image after rasterization is compared with the HR ground truth to calculate $L_{1}$ loss. During back propagation, the gradient will pass through the differentiable rasterization to optimize the model backbone, guiding the framework to convert the input LR image into the corresponding image-conditioned 2D Gaussians $\mathcal{G}$ in a learning-based manner.
4 Experiments
4.1 Experimental Setup
Implementation Details. We utilize the widely-used DIV2K [41] dataset as training set. We follow [10, 2] to fix LR patches to $48\times 48$. The scaling factor $s$ is randomly selected within the range $[1.0,4.0]$. We first crop $48s\times 48s$ GT patch from the original full-size GT image, and then apply bicubic down-sampling [2] to obtain the corresponding LR patches. Following LIIF [10], we employ EDSR [32] and RDN [64] as the image encoder backbones to extract features from the input LR images. We train our GSARS on $4$ NVIDIA A100 GPUs for $500,000$ iterations with batch size 64. The initial learning rate is $2e^{-4}$, and halves at $250,000$, $400,000$, $450,000$, $475,000$ iterations. The Adam [25] optimizer is utilized. To speed up the convergence of the training progress, we set the number of warm-up iterations to $2,000$. The window size $k$ in Condition injection block and Gaussian Interaction Block is set to $12$, we set the proportion parameter $m=16$, the dimension of Gaussian embedding $d=180$, and stack $L=6$ Gaussian interaction blocks.
Comparison Methods. We compare our proposed GSASR with Meta-SR [20], LIIF [10], LTE [27], SRNO [49], LINF [54], CiaoSR [2], LMF [18] and GaussianSR [19].
Evaluation Protocols. We evaluate the competing models on Set5 [1], Set14 [57], DIV2K100 [41], Urban100 [21], BSDS100 [35], Manga109 [36], General100 [15], and LSDIR [28] datasets under scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$. To validate the generalization on higher scaling factors, we further report the results with ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$ scaling factors on DIV2K100 [41] and LSDIR [28] datasets. The PSNR, SSIM [48], LPIPS [60] and DISTS [13] indices are used for performance evaluation. Note that some existing methods [10, 2] calculate PSNR on RGB channels for DIV2K [41], but on the Y channel of YCbCr space for other datasets. To unify the testing protocols, we calculate PSNR/SSIM on the Y channel across all datasets and all competing methods for comparison. For a fair comparison, we download all competing models from their official websites, utilize the same data to generate SR results and the same evaluation codes to compute the metrics.
For the comparison of computational cost, we crop $100$ GT patches with $720\times 720$ resolution from DIV2K [41], and use bicubic interpolation to generate the corresponding LR images with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$. We report the average inference time (ms) and GPU memory usage (MB) on a single NVIDIA A100 GPU. The computational cost is calculated over the full SR process, including encoder, GS decoder and rendering.
@@ 4.2 Experiment Results
Quantitative Results. Table 1 shows the numerical results of our proposed GSASR and other comparison methods with EDSR-baseline [32] encoder. One could see that GSASR outperforms existing methods in terms of both fidelity metrics (PSNR and SSIM [48]) and perceptual quality metrics (LPIPS [60] and DISTS [13]) under scaling factors from ×$\times 2$ to ×$\times 30$. It obtains significant improvements over the methods based on INR, exhibiting greater potential for ASR. Compared with the recent GaussianSR method [19], which also employs GS for feature representation, GSASR shows clear advantages. GSASR predicts the position of Gaussians in an end-to-end manner, while GaussianSR [19] simply replaces the position parameters with the fixed position of the RGB pixels of the LR image and lacks sufficient representation capability. Besides, GaussianSR lacks GPU/CUDA-based rasterization to accelerate the inference speed and it fixes the number of Gaussians to $100$, no matter how large the scaling factor is. Therefore, the performance of GaussianSR drops a lot under high magnification factors. In contrast, we present an efficient 2D GPU/CUDA-based rasterization to embed more Gaussians to promote ASR performance. As a result, although GSASR is trained with a scaling factor from $1.0$ to $4.0$, it could not only obtain excellent performance for in-distribution scaling factors (i.e., scaling factor less than $4$), but also show strong capability in dealing with out-of-distribution situations (i.e., scaling factor larger than $4$).
Due to the limited pages, more results on Set5 [1], Set14 [57], Urban100 [21], BSDS100 [35], Manga109 [36], General100 [15] with EDSR-baseline [32] and RDN [64] encoder could be found in the supplementary file.
@@
Qualitative Results. Fig. 4 and Fig. 5 show the qualitative comparison of competing methods. We can see that the compared methods mostly produce blurry details (e.g., the holes in Fig. 4), while GSASR generates much clearer edges and details, owing to the strong representation ability of 2D Gaussians. More visualization results can be found in the supplementary file. To explain why GSASR works well for the ASR task, we visualize the positions together with the shape of Gaussians in Fig. 6. One could see that the positions of Gaussians tend to cluster in areas with complex textures (such as the windows), while they are uniformly distributed in flat areas (such as the sky). This validates that our proposed strategy is highly suitable for restoring complex texture details. Meanwhile, from the blue and green boxes in Fig. 6, one could find that the Gaussians will adaptively adjust their orientations and scales based on the direction and shape of the textures.
@@
Computational Costs. Table 2 shows that GSASR surpasses all competing methods in most of the fidelity (PSNR, SSIM) and perceptual (LPIPS, DISTS) measures. In terms of speed, GSASR is nearly twice as fast as the state-of-the-art method CiaoSR [2] under ×3,×4$\times$ 3$\times$ 4× 3 , × 4 scaling factors. The speed advantage becomes more significant when handling with larger scaling factors. Thanks to the efficient 2D GPU/CUDA-based rasterization, GSASR could render an image with high speed without sacrificing the reconstruction fidelity. The computational cost with RDN backbone can be found in the supplementary file.
@@
Ablation Study. We conduct ablation studies on (1) the number of Gaussians $N$, (2) the functionality of the reference position, (3) the rasterization ratio $r$, (4) the dimension $d$ of Gaussian embedding, abd (5) the window size $k$ in condition injection block and Gaussian Interaction block. Details are presented in the supplementary file.